# Guided Bayesian Optimization: Data-Efficient Controller Tuning with Digital Twin

### v1

@@ VI-C Controller Tuning Time
In each experiment on the physical systems, starting from an initial condition, we measure the step response during $10$ s and $0.30$ s for the rotary and linear motor systems, respectively. To decrease the noise level, we repeat it twice per experiment for the DC motor setup to take the average estimated value per each performance metric.
Each batch of experiments consists of a single experiment conducted for the offline initial data set and specific experiments per each data-driven tuning, and on average, $121$ experiments on DT in the guided BO method.
DT does not require a sophisticated model that would be time-consuming to calculate, and each BO iteration on the DT requires only $0.09$ s of overdue computation time. So, the experiments on the physical system demand longer than the DT iterations.
Furthermore, for the PGM nominal controller, one needs to identify the plant model that, in our case, takes $7.66$ s for the linear system and $40.35$ s for the rotary system using the Labview-based real-time tool introduced in [43]. PGM method also takes $2$ ms to calculate the controller parameters given the identified LTI model.
@@ VII Conclusion
In this paper, we guided the data-driven Bayesian optimization-based controller with a digital twin according to the uncertainty level of the GP model. The digital twin is approximated with available data during the system’s operation without additional experiments. We demonstrated that our guided BO approach considerably improves the controller tuning data efficiency, generalizes across the system and industrial controller structures, and converges faster to optimum values. Namely, we proved the guided BO’s superior performance on a noisy linear servo motor and DC rotary motor real-time hardware.
@@ VIII Acknowledgement
We thank Matthias Geissmann for his support in providing the LinMot equipment necessary to experimentally validate the approach and in discussions about the linear motor properties.

### v2

@@ VI-C Controller Tuning Time
In each experiment on the physical systems, starting from an initial condition, we measure the step response during $10$ s and $0.30$ s for the rotary and linear motor systems, respectively. To decrease the noise level, we repeat it twice per experiment for the DC motor setup and take the average estimated value per each performance metric.
Each batch of experiments consists of a single experiment conducted for the offline initial data set and specific experiments per each data-driven tuning, and on average, $121$ experiments on DT in the guided BO method. The DT does not require a sophisticated model that would be time-consuming to calculate, and each BO iteration on the DT requires only $0.09$ s of overdue computation time. So, the experiments on the physical system demand longer than the DT iterations.
For the PGM nominal controller, one needs to identify the plant model, which takes $7.66$ s for the linear system and $40.35$ s for the rotary system using the Labview-based real-time tool introduced in [51]. The PGM method takes $2$ ms to calculate the controller parameters given the identified LTI model.
@@ VII Conclusion
In this paper, we guided the data-driven Bayesian optimization-based controller with a digital twin according to the uncertainty level of the GP model. The digital twin is approximated with available data during the system’s operation without additional experiments. As long as the digital twin plant roughly captures the overall behavior of the system far away from the optimum region, it is going to be helpful to guide the Bayesian optimizer. We showed that the digital twin carries out the exploration duty of the optimizer whereas the exploitation is performed on the real system. We demonstrated that our guided BO approach considerably improves the controller tuning data efficiency, generalizes across the system and industrial controller structures, and converges faster to optimum values. Namely, we proved the guided BO’s superior performance on a noisy linear servo motor and DC rotary motor real-time hardware.
@@ VIII Acknowledgement
We thank Matthias Geissmann for providing the LinMot equipment and useful discussions about its properties.